Shift report 20030317 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Mar 17 2003 - 16:11:22 EST

  • Next message: Apache: "Shift report 20030318 16:00-24:00"
    ------------------------------------------------------------
    Date: Mon Mar 17 16:11:22 2003, Shift: 08:00-16:00
    Supervisor: sjs, with: rl
    Log book#14 pages 66 - 67
    
    Shift Summary:
    ------------------------------------------------------------
    Air conditioning problem kept us off-line all morning. 
    >From 8:00-11:00 there was a good store that we missed.
    Air conditioner fixed around 11:30, but various RHIC 
    problems resulted in no additional stores during the day.
    We have a problem with the magnets tripping off...
    They now don't seem to stay on for more than an hour or so.
    MCR has been informed.
    A new store is being prepared.  Hopefully we will
    be able to take advantage of it.
    Run 8470 is tpc pedestal. Run 8471 is dvm.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon Mar 17 2003 - 16:16:06 EST